# StyleCity: Large-Scale 3D Urban Scenes Stylization with Vision-and-Text Reference via Progressive Optimization

### v1

@@ Abstract
Creating large-scale virtual urban scenes with variant styles is inherently challenging. To facilitate prototypes of virtual production and bypass the need for complex materials and lighting setups, we introduce the first vision-and-text-driven texture stylization system for large-scale urban scenes, StyleCity. Taking an image and text as references, StyleCity stylizes a 3D textured mesh of a large-scale urban scene in a semantics-aware fashion and generates a harmonic omnidirectional sky background. To achieve that, we propose to stylize a neural texture field by transferring 2D vision-and-text priors to 3D globally and locally. During 3D stylization, we progressively scale the planned training views of the input 3D scene at different levels in order to preserve high-quality scene content. We then optimize the scene style globally by adapting the scale of the style image with the scale of the training views. Moreover, we enhance local semantics consistency by the semantics-aware style loss which is crucial for photo-realistic stylization. Besides texture stylization, we further adopt a generative diffusion model to synthesize a style-consistent omnidirectional sky image, which offers a more immersive atmosphere and assists the semantic stylization process. The stylized neural texture field can be baked into an arbitrary-resolution texture, enabling seamless integration into conventional rendering pipelines and significantly easing the virtual production prototyping process. Extensive experiments demonstrate our stylized scenes’ superiority in qualitative and quantitative performance and user preferences.
@@
3D Stylization City Stylization Photorealistic Style Transfer Neural Style Transfer
@@ 1 Introduction
Large-scale urban digital twins, reconstructed from satellite and UAV imagery, are widely utilized in various domains such as video games, 3D city visualization, and virtual reality applications. In virtual production, the demand for visually appealing and unique representations of reconstructed meshes is high. Yet, simulating the visual appearance of virtual scenes—like the time simulation in Google Earth—often fails to produce truly impressive visuals. This shortfall is typically due to the extensive need for laborious material design and meticulous illumination setups for each instance. To address this challenge, we propose an automated system for textured mesh stylization in this work. It serves as a lightweight alternative for the re-styling of virtual scene appearances, utilizing 3D neural style transfer techniques.
@@
To narrow the gap, we propose to use a UV-based neural texture field to model the scene appearance. The UV-based neural texture field is composed of a multi-resolution feature grid for UV encoding and a multilayer perceptron (MLP) for feature decoding. As a compact hybrid representation, it reduces resource usage while maintaining global style consistency to some extent. After initializing the neural texture field with the original UV texture map, we perform optimization by minimizing global content and style losses among multiple planned views.
@@ 2 Related Work
In early works for 3D stylization, researchers applied interactive and iterative optimization on 3D data such as mesh texture and geometry to achieve non-photorealistic stylization [12, 53, 16, 33]. Later, some methods [59, 5] trained generative neural networks with large 3D data and facilitated 3D appearance and geometry style transfer. For example, 3DStyleNet [59] transferred both shape and texture between 3D models with a textured mesh exemplar by a 3D generative model. 3D data collection is tedious and time-consuming. Fortunately, empowered by large-data pre-trained 2D neural models such as VGG, CLIP [45], and Stable Diffusion [48], there appeared a lot of automatic vision or text example-based 2D neural style transfer approaches [14, 21, 35, 42, 13, 29, 27, 58]. Inspired by this, researchers tend to lift these 2D priors to 3D representations and achieve automatic 2D example-based 3D stylization. In the following sections, we review 3D neural style transfer methods on meshes and neural field representations, which transfer style with vision or text guidance and preserve to some degree original structure and geometry. We refer to [7] for more 3D stylization works on other 3D data.
2.1 Mesh-based Stylization
@@
Kato et al. [25] proposed a neural renderer for mesh such that it enables artistic geometric and appearance stylization for the object mesh using image-example-based neural style transfer technique [14]. When it turns to image-guided 3D indoor scene artistic stylization, StyleMesh [18] introduced a texture optimization scheme specifically designed for 3D reconstructed indoor meshes, incorporating depth and angle information. Later, with the emergence of large-language model (LLM), text-guided stylization on mesh geometry and appearance [37, 36], text-guided texture transfer or synthesis [24, 31, 47, 6, 57] became a popular topic in academia. [24] accomplished 3D semantic style transfer for indoor textures with text guidance. [37, 36] achieved text-guided object geometric and texture stylization using a vertex-based neural style field network. They largely rely on fine-grained triangulation on the mesh and cannot synthesize detailed patterns or motifs. [31, 57] took object materials into account, achieving text-guided texture stylization with material properties. With the advent of large-data pre-trained diffusion models [48], recent works [47, 6, 56, 4, 28, 52] proposed to use the pre-trained diffusion generative models [61, 48, 3] to better supervise text-guided photorealistic texture synthesis for objects or rooms. [57] utilized 2D diffusion priors [44] for object texture stylization. However, globally optimizing explicit texture requires huge computational resources, i.e., GPU memory. These methods generally focus on small-scale scenes and lack scalability.
2.2 Neural Field Stylization
@@
There appeared numerous neural-field-based stylization works for implicit reconstructed scenes [19, 10, 22, 11, 40, 64, 55, 15]. Some data-driven generative training models [19, 34] support zero-shot stylization for a scene, but require large training references and sophisticated model design. By contrast, feature optimized-based approaches usually have more impressive results without other style dilutions for one model [60, 63, 41]. For example, ARF [60] proposed a nearest-neighbor feature matching loss for local style matches in latent space transferring local artistic patterns. Empowered by 2D generative models, there are generative-model supervised optimization methods, such as InstructN2N [15], which utilized an alternative multi-view updating scheme to progressively optimize the 3D scene. However, multi-view supervision by independent one-shot generated images easily tends to collapse due to 3D consistency. These neural field stylization works utilized 3D neural field representations to maintain global style consistency and render novel views via volume rendering. However, it is not wise to convert mesh-based rendering into volume rendering for mesh stylization which is prone to generating blurry views.
@@ 3.1 Data Pre-processing
3.1.1 Pivot-Based View Planning.
Naive scene center view planning methods which are used to stylize single objects [37, 47, 26] are insufficient for large-scale surface coverage. Direct application results in blurry texture with broken structure after stylizaiton as shown in Fig. 2(b). We developed a pivot-based view planning method that overcome their limitations. We uniformly sample $P$ camera positions on the upper and side faces of the mesh bounding box with a proper offset. The mesh is then subdivided into $r$ sub-regions, and the centroids of these regions are used as camera viewing points. Finally, we obtain $P\times r$ pivot views, locating at $P$ pivot positions and looking at $r$ centroids. Fig. 2(a) shows a pivot view planning example. Pivot views cover the majority of visible surfaces with associated semantics, serving as initialization for the novel views (Fig. 2(a) green camera) augmenting training, referred to Sec. 3.3.1 Multi-Scale Progressive Optimization.
@@
3.1.2 Urban Scene Segmentation.
For semantics-aware urban scenes stylization, we particularly consider classes of interest in architectural scenarios for segmentation, including "sky", "building", "window", "road", "person", "plant", "car", "water" and "lights". We adapted and fine-tuned Mask2Former [9], a pre-trained 2D semantic segmentation model, and implemented an automated 2D-to-3D segmentation tool via texture mapping for input 3D models. We also segment the input style reference image using the same segmentation model.
@@ 3.2 Neural Texture Field
Neural Rendering. Given a camera pose, we rasterize mesh and retrieve UVs to query texture model $T_{\Theta}(\cdot)$. We then obtain corresponding texture RGB values, and then resemble values into a rendered image.
@@ 3.3 Content and Style Joint Progressive Optimization
During stylization, we utilize content features from rendered views and Laplacian values for supervision, aiming to preserve texture content and scene identity. To maintain content integrity and photorealism, we mask out the background areas during the training phase.
@@
$\mathcal{L}_{c}=\sum_{\mathit{l\in l_{c}}}\|\mathit{F}^{l}(c)-\mathit{F}^{l}(z)\|^{2}_{2},\vspace{-0.3cm}$ (1)
@@
$\mathcal{L}_{pht}=\sum_{h=1}^{3}(\mathcal{V}^{z}_{h})^{T}\mathcal{M}^{c}\mathcal{V}^{z}_{h},\vspace{-0.3cm}$ (2)
@@
$\begin{split}\arg\min_{\hat{s}\in S}\sum&\|\mathcal{D}(c)-\mathcal{D}(\hat{s})\|^{2}_{2},\end{split}\vspace{-0.2cm}$ (3)
@@
$\mathcal{L}_{gs}=\sum_{l\in l_{s}}(\|\mu(\mathit{{F}}^{l}(\hat{s}))-\mu(\mathit{F}^{l}(z))\|^{2}_{2}+\|\sigma(\mathit{{F}}^{l}(\hat{s}))-\sigma(\mathit{F}^{l}(z))\|^{2}_{2}),\vspace{-0.5em}$ (4)
where $\hat{s},z$ respectively denote the matched style image and output stylized image; feature layers $l_{s}=\{2,3,4\}$; $\mu$ and $\sigma$ indicate mean and standard deviations. To transfer textual style features for holistic style semantics harmony, we apply global and directional CLIP losses [45, 13]:
@@
Semantic Hierarchical Re-match. During style transfer, it may occur mismatched and missed semantics between content view and style image as shown in Fig. 6. Content scene contains areas in semantics "building", "window", "plant", "water" and "road" to be stylized, while style image only offers reference semantics in "sky", "building", "window" and "light". To obtain references for missed semantics (e.g., "plant" and "water"), we designed a hierarchy graph for semantics re-matching based on illumination effects in the real world. For example, cars, trees, and people on the road should share similar illumination of the road; all elements have the same ambient illumination effect from the sky, etc. In Fig. 6, "water" content can match "sky" style luminance, "plant" content can match "road" reference luminance. If the matched style patch fails to provide style of some classes, we turn back to full-scale style reference for help.
@@
We now have the texture stylization model with all the above-mentioned losses:
$\vspace{-0.3cm}\begin{split}\mathcal{L}_{total}&=\underbrace{\lambda_{c}\mathcal{L}_{c}+\lambda_{pht}\mathcal{L}_{pht}}_{\mathcal{L}_{content}}+\underbrace{\lambda_{gs}\mathcal{L}_{gs}+\lambda_{ts}\mathcal{L}_{clip}+\lambda_{ls}\mathcal{L}_{ls}}_{\mathcal{L}_{style}},\end{split}\vspace{-0.3cm}$ (8)
@@ 4 Experiments
We conducted experiments on UV textured models in different cities collected by Google 3D Tiles API [2]. For evaluation, we rendered a sequence of testing views around each model. For 2D style images, we referred to city-view evaluation images in [8, 35], and applied the text-guided synthesis model [61] to generate some style images given customized style text prompts. Source texts follow templates such as "a photo in the day", and non-customized target texts are generated by [43] given style images. All experiments were conducted on NVIDIA RTX 3090 GPU. We implemented our method based on Pytorch and Pytorch3D [46], and used tiny-cuda-nn [38] for multi-resolution grid-based hash encoding and network training acceleration. Resolution $512\times 512$ of views was used for optimization.
@@
We compared our method with the image-guided explicit texture stylization StyleMesh [18], image- and text-guided NeRF stylization methods ARF [60] and instructN2N [15]. We ran baselines using official source codes and increased the content loss hyperparameter used in StyleMesh and ARF by 10 times for photorealism. We took the same pivot views (Sec 3.1.1) as their training input. We evaluated the performance of foreground content similarity by edge-SSIM (eSSIM) and masked LPIPS [62] between source and stylized testing views, and semantic style similarity by image-text CLIP score [17] between the text reference and stylized testing views with the same sky background. Moreover, we conducted a user study on content preservation and style matching of stylized models. Table 1 and Fig. 8 demonstrate quantitative and qualitative comparisons.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption

### v2

# Corresponding author

@@ Abstract
Creating large-scale virtual urban scenes with variant styles is inherently challenging. To facilitate prototypes of virtual production and bypass the need for complex materials and lighting setups, we introduce the first vision-and-text-driven texture stylization system for large-scale urban scenes, StyleCity. Taking an image and text as references, StyleCity stylizes a 3D textured mesh of a large-scale urban scene in a semantics-aware fashion and generates a harmonic omnidirectional sky background. To achieve that, we propose to stylize a neural texture field by transferring 2D vision-and-text priors to 3D globally and locally. During 3D stylization, we progressively scale the planned training views of the input 3D scene at different levels in order to preserve high-quality scene content. We then optimize the scene style globally by adapting the scale of the style image with the scale of the training views. Moreover, we enhance local semantics consistency by the semantics-aware style loss which is crucial for photo-realistic stylization. Besides texture stylization, we further adopt a generative diffusion model to synthesize a style-consistent omnidirectional sky image, which offers a more immersive atmosphere and assists the semantic stylization process. The stylized neural texture field can be baked into an arbitrary-resolution texture, enabling seamless integration into conventional rendering pipelines and significantly easing the virtual production prototyping process. Extensive experiments demonstrate our stylized scenes’ superiority in qualitative and quantitative performance and user preferences. Project page: https://chenyingshu.github.io/stylecity3d.
@@
3D Stylization City Stylization Neural Style Transfer
@@ 1 Introduction
Large-scale urban digital twins, reconstructed from satellite and UAV imagery, are widely utilized in various domains such as video games, 3D city visualization, and virtual reality applications. In virtual production, the demand for visually appealing and unique representations of reconstructed meshes is high. Yet, simulating the visual appearance of virtual scenes—like the time simulation in Google Earth—often fails to produce truly impressive visuals. This shortfall is typically due to the extensive need for laborious material design and meticulous illumination setups for each instance. To achieve automatically re-styling scene appearances, we seek to utilize 3D neural style transfer techniques.
@@
To narrow the gap, we propose to use a UV-based neural texture field to model scene appearances. The UV-based neural texture field is composed of a multi-resolution feature grid for UV encoding and a multilayer perceptron (MLP) for feature decoding. As a compact hybrid representation, it reduces resource usage while maintaining global style consistency to some extent. After initializing the neural texture field with the original texture map, we perform optimization by minimizing global content and style losses among multiple planned views.
@@ 2 Related Work
In early works for 3D stylization, researchers applied interactive and iterative optimization on 3D data such as mesh texture and geometry to achieve non-photorealistic stylization [12, 53, 16, 33]. Later, some methods [59, 5] trained generative neural networks with large 3D data and facilitated 3D appearance and geometry style transfer. For example, 3DStyleNet [59] transferred both shape and texture between 3D models with a textured mesh exemplar by a 3D generative model. 3D data collection is tedious and time-consuming. Fortunately, empowered by large-data pre-trained 2D neural models such as VGG, CLIP [45], and Stable Diffusion [48], there appeared a lot of automatic vision or text example-based 2D neural style transfer approaches [14, 21, 35, 42, 13, 29, 27, 58]. Inspired by this, researchers tend to lift these 2D priors to 3D representations for automatic 2D example-based 3D stylization. Next, we review 3D neural style transfer methods on meshes and neural field representations, which transfer style with vision or text guidance and preserve to some degree original structure and geometry. We refer to [7] for more 3D stylization works on other 3D data.
@@
Mesh-based Stylization. Kato et al. [25] proposed a neural renderer for mesh such that it enables artistic geometric and appearance stylization for the object mesh using image-example-based neural style transfer technique [14]. When it turns to image-guided 3D indoor scene artistic stylization, StyleMesh [18] introduced a texture optimization scheme specifically designed for 3D reconstructed indoor meshes, incorporating depth and angle information. Later, with the emergence of large-language model (LLM), text-guided stylization on mesh geometry and appearance [37, 36], text-guided texture transfer or synthesis [24, 31, 47, 6, 57] became a popular topic in academia. [24] accomplished 3D semantic style transfer for indoor textures with text guidance. [37, 36] achieved text-guided object geometric and texture stylization using a vertex-based neural style field network. They largely rely on fine-grained triangulation on the mesh and cannot synthesize detailed patterns or motifs. [31, 57] took object materials into account, achieving text-guided texture stylization with material properties. With the advent of large-data pre-trained diffusion models [48], recent works [47, 6, 56, 4, 28, 52] proposed to use the pre-trained diffusion models [61, 48, 3] to better supervise text-guided photorealistic texture synthesis for objects or rooms. [57] utilized 2D diffusion priors [44] for object texture stylization. However, globally optimizing explicit texture requires huge computational resources, i.e., GPU memory. These methods generally focus on small-scale scenes and lack scalability.
@@
Neural Field Stylization. There appeared numerous neural-field-based stylization works for implicit reconstructed scenes [19, 10, 22, 11, 40, 64, 55, 15]. Some data-driven generative training models [19, 34] support zero-shot stylization for a scene, but require large training references and sophisticated model design. By contrast, feature optimized-based approaches usually have more impressive results without other style dilutions for one model [60, 63, 41]. For example, ARF [60] proposed a nearest-neighbor feature matching loss for local style matches in latent space transferring local artistic patterns. Empowered by 2D generative models, there are generative-model supervised optimization methods, such as InstructN2N [15], which utilized an alternative multi-view updating scheme to progressively optimize the 3D scene. However, multi-view supervision by independent one-shot generated images easily tends to collapse due to 3D inconsistency. These neural field stylization works utilized 3D neural field representations to maintain global style consistency and render novel views via volume rendering. However, it is not wise to convert mesh-based rendering into volume rendering for mesh stylization which is prone to generating blurry views.
@@ 3.1 Data Pre-processing
Pivot-Based View Planning. Naive scene center view planning methods applied to stylize single objects [37, 47, 26] are insufficient for large-scale surface coverage. Direct application results in blurry texture with broken structure after stylization as shown in Fig. 2(b). We developed a pivot-based view planning method that overcome their limitations. We uniformly sample $P$ camera positions on the upper and side faces of the mesh bounding box with a proper offset. The mesh is then subdivided into $r$ sub-regions, and the centroids of these regions are used as camera viewing points. Finally, we obtain $P\times r$ pivot views, locating at $P$ pivot positions and looking at $r$ centroids. Fig. 2(a) shows a pivot view planning example. Pivot views cover the majority of visible surfaces with associated semantics, serving as initialization for the novel views (Fig. 2(a) green camera) augmenting training, referred to Sec. 3.3.1 Multi-Scale Progressive Optimization.
@@
Urban Scene Segmentation. For semantics-aware urban scenes stylization, we particularly consider classes of interest in architectural scenarios for segmentation, including "sky", "building", "window", "road", "person", "plant", "car", "water" and "lights". We adapted and fine-tuned Mask2Former [9], a pre-trained 2D semantic segmentation model, and implemented an automated 2D-to-3D segmentation tool via texture mapping for input 3D models. We also segment the input style reference image using the same segmentation model.
@@ 3.2 Neural Texture Field
Neural Rendering. Given a camera pose, we rasterize mesh and retrieve UVs to query texture model $T_{\Theta}(\cdot)$. Accordingly, we then obtain corresponding texture RGB values, and then resemble values into a rendered image.
@@ 3.3 Content and Style Joint Progressive Optimization
During stylization, we utilize not only content features but also Laplacian values of rendered views for supervision, preserving texture content and scene identity. During training, we mask out sky background areas.
@@
$\mathcal{L}_{c}=\sum_{\mathit{l\in l_{c}}}\|\mathit{F}^{l}(c)-\mathit{F}^{l}(z)\|^{2}_{2},$ (1)
@@
$\mathcal{L}_{pht}=\sum_{h=1}^{3}(\mathcal{V}^{z}_{h})^{T}\mathcal{M}^{c}\mathcal{V}^{z}_{h},$ (2)
@@
$\begin{split}\arg\min_{\hat{s}\in S}\sum&\|\mathcal{D}(c)-\mathcal{D}(\hat{s})\|^{2}_{2},\end{split}$ (3)
@@
$\mathcal{L}_{gs}=\sum_{l\in l_{s}}(\|\mu(\mathit{{F}}^{l}(\hat{s}))-\mu(\mathit{F}^{l}(z))\|^{2}_{2}+\|\sigma(\mathit{{F}}^{l}(\hat{s}))-\sigma(\mathit{F}^{l}(z))\|^{2}_{2}),$ (4)
where $\hat{s},z$ respectively denote the matched style image and output stylized image; feature layers $l_{s}=\{2,3,4\}$; $\mu$ and $\sigma$ indicate mean and standard deviations.
To transfer textual style features for holistic style semantics harmony, we apply global and directional CLIP losses [45, 13]:
@@
Semantic Hierarchical Re-match. During style transfer, it may occur mismatched and missed semantics between content view and style image as shown in Fig. 6. Content scene contains semantics "building", "window", "plant", "water" and "road", while style image lacks corresponding semantics such as "plant" and "water". To obtain references for missed semantics, we designed a hierarchy graph for semantics re-matching based on illumination effects in the real world. For example, cars, trees, and people on the road should share similar illumination of the road; all elements have the same ambient illumination from the sky, etc. In Fig. 6, "water" content can match "sky" style luminance, "plant" content can match "road" reference luminance. If the matched style patch fails to provide style of some classes, we turn back to full-scale style reference for help.
@@
We now have the texture stylization model with all losses:
$\begin{split}\mathcal{L}_{total}&=\underbrace{\lambda_{c}\mathcal{L}_{c}+\lambda_{pht}\mathcal{L}_{pht}}_{\mathcal{L}_{content}}+\underbrace{\lambda_{gs}\mathcal{L}_{gs}+\lambda_{ts}\mathcal{L}_{clip}+\lambda_{ls}\mathcal{L}_{ls}}_{\mathcal{L}_{style}},\end{split}$ (8)
@@ 4 Experiments
We conducted experiments on UV textured models in different cities collected by Google 3D Tiles API [2]. For evaluation, we rendered a sequence of testing views around each model. For 2D style images, we referred to city-view evaluation images in [8, 35], and applied the text-guided synthesis model [61] to generate some style images given customized style text prompts. Source texts follow templates such as "a photo in the day", and non-customized target texts are generated by [43] given style images. All experiments were conducted on RTX 3090 GPU. We implemented our method based on Pytorch and Pytorch3D [46], and used tiny-cuda-nn [38] for multi-resolution grid-based hash encoding and network training acceleration. Resolution $512\times 512$ was used for optimization.
@@
Given the downgraded performance of image or video stylization in 3D scenarios regarding 3D geometric and appearance consistency [10, 19, 22], our focus lies on 3D stylization baselines. We compared our method with the image-guided explicit texture stylization StyleMesh [18], image- and text-guided NeRF stylization methods ARF [60] and InstructN2N [15]. We ran baselines using official source codes and increased the content loss hyperparameter used in StyleMesh and ARF by 10 times for photorealism. We took the same pivot views (Sec. 3.1) as their training input. We evaluated the performance of foreground content similarity by edge-SSIM (eSSIM) and masked LPIPS [62] between source and stylized testing views, and semantic style similarity by image-text CLIP score [17] between the text reference and stylized testing views with the same sky background. Moreover, we conducted a user study on content preservation and style matching of stylized models. Table 1 and Fig. 8 demonstrate quantitative and qualitative comparisons.